Show that the two spheres x²+y²+z²+6y+2z+8=0 and x²+y²+z²+6x+8y+4z+20=0

 are orthogonal.


1
Expert's answer
2020-09-14T19:16:18-0400

"2*u_1*u_2+2*v_1*v_2+2*w_1*w_2=d_1+d_2" --- it is a orthogonality condition

"x^2+y^2+z^2+6*y+2*z+8=0" --- the first sphere

"u_1 = 0\/2 = 0" ; "v_1 = 6\/2 = 3" ; "w_1 = 2\/2 = 1" ; "d_1 = 8"

"x^2 + y^2+z^2+6*x+8*y+4*z+20 = 0" --- the second sphere

"u_2 = 6\/2 = 3" ; "v_2 = 8\/2 = 4" ; "w_2 = 4\/2 = 2" ; "d_2 = 20"

As the first sphere is orthogonal to the second sphere, we must have

"2*0*3+2*3*4+2*2*1=8+20"

"24+4=28"

"28=28" - equality is true

so these two spheres are orthogonal